OpenStudy (anonymous):

Write an expression for the following real-world situation. If Mr. Jackson earns $1000 for every used car he sells and $2000 for every new car he sells, how much would he make if he sold "u" used cars and "n" new cars? Write a real-world situation of your own and a VARIABLE expression to go with it. Help ?

OpenStudy (anonymous):

if he sells u used cars he will make \[u \times $1000\] and if he sells n new cars he will make \[n \times $2000\]

OpenStudy (anonymous):

So it would be u X 1000 + n X 2000 ?

OpenStudy (anonymous):

yeah if he sells both in that case \[u \times $1000\] + \[n \times $2000\]

OpenStudy (anonymous):

OKie dokieee thanks (: